# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

### v2

@@ Abstract
Despite significant advancements in text generation and reasoning, Large Language Models (LLMs) still face challenges in accurately performing complex arithmetic operations. To achieve accurate calculations, language model systems often enable LLMs to generate code for arithmetic operations. However, this approach compromises speed and security and, if finetuning is involved, risks the language model losing prior capabilities. We propose a framework that enables exact arithmetic in a single autoregressive step, providing faster, more secure, and more interpretable LLM systems with arithmetic capabilities. We use the hidden states of an LLM to control a symbolic architecture which performs arithmetic. Our implementation using Llama 3 8B Instruct with OccamNet as a symbolic model (OccamLlama) achieves 100% accuracy on single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), outperforming GPT 4o and on par with GPT 4o using a code interpreter. OccamLlama also outperforms GPT 4o both with and without a code interpreter on mathematical problem solving benchmarks involving challenging arithmetic, thus enabling small LLMs to match the arithmetic performance of even much larger models. We will make our code public shortly.
@@ 1 Introduction
    We benchmark on mathematical problem solving tasks, showing that OccamLlama can sustain long generations. OccamLlama outperforms both GPT 4o and GPT 4o with code interpreter on benchmarks involving challenging arithmetic.
@@ 2 Related Work
Table 1: OccamLLM is the only approach to improving the arithmetic capabilities of a pretrained LLM which 1) enables single pass arithmetic, 2) does not risk catastrophic forgetting from finetuning, 3) does not require arbitrary code execution, and 4) provides an interpretable process.
@@ 3.1 OccamLLM: Combining a Language Model with a Symbolic Model
In the subsections below, we describe the OccamLLM system which from our experiments we find to be most performant, even oupterforming GPT 4o in several benchmarks. For an analysis of alternate architectures and losses, see Appendix C.
@@ 3.1.1 OccamNet
OccamNet is a symbolic architecture that provides an interpretable way of parametrizing probability distributions over a space of functions [11]. We leave a more thorough explanation of OccamNet to [11] and Appendix D, describing only the relevant components here.
@@ 3.1.2 OccamLLM Decoder
where $\mathbf{h}_{j}$ are the hidden states of the $j$th layer of the language model, $w_{i,j}$ are trainable weights, $\text{MLP}_{i}$ are two-layer multilayer perceptrons (MLPs), and $\mathbf{W}^{*(i)}$ are untrained weights which initialize all functions to have approximately equal probabilities according to the initialization scheme described in [11] and explained in Appendix D.4.
@@ 3.2.1 OccamNet Decoder Training Data
For training the decoder that controls the weights of OccamNet, we created two types of examples, single queries and concatenated queries. For single queries, we select a single prompt from the problems generated as discussed in Section 3.2. We use the Llama 3 8B Instruct chat template and fill in the query as the user input and the result as the assistant response, prepending “Answer = ” to the later in randomly selected samples (see Appendix A.1 for further details). For the concatenated queries of examples, we select a random number of prompts and concatenate the query-response pairs without using the Llama 3 8B Instruct chat template. The OccamNet decoder is trained to predict only the results of the last query in the sequence. This strategy helps OccamLLM to learn which operation to perform without becoming confused by earlier text, which is useful for continuous generation. To create the training dataset, each example is sampled by first randomly selecting whether to create a single or concatenated query, then randomly selecting the type(s) of prompt(s) used, and finally randomly sampling the input values from the range corresponding to each selected prompt.
@@ 3.2.2 OccamLLM Switch Training Data
To train the switch, we generate examples of possible LLM outputs for given input expressions, and label the outputs with sequences of 0s or 1s corresponding to whether the language model or the OccamNet output should be used for the next token. Some examples correspond to the prompts described in Section 3.2. For such examples, the LLM output is set to “The answer is” or “Answer = ” and the label sequence is all 0s with a 1 at the last token to indicate the system should use OccamNet only to compute the answer. We also manually created and labeled several other examples for diverse scenarios to explicitly teach the system in which cases it should or should not use OccamNet.
To create the training dataset, we concatenate a random number of the above user input - assistant output pairs in a conversational fashion, using the Llama 3 8B Instruct chat template.
@@ 3.3 OccamLLM Training
We train the OccamLLM decoder and the switch separately, as they do not share weights. In all cases, the weights of the LLM are kept frozen. In the first step, we train the system to predict the answer to examples generated by the method explained in Section 3.2.1. The OccamNet decoder processes the hidden states corresponding to the last token of the response and sets the weights of OccamNet such that the correct arithmetic expression is sampled. In this step, we use a rescaled REINFORCE [34] loss, which can also be interpreted as a Monte-Carlo estimate of the cross-entropy loss (see Appendix C.2):
@@
Minimizing this loss steers the decoder towards assigning higher probabilities to the functions that maximize the reward $R(f(x),y)$, which measures the similarity between the correct answer $y$ and the prediction of OccamNet $f(x)$. We find setting $R(f(x),y)=1$ if $f(x)=y$, and 0 otherwise, most effective. We discuss the OccamNet loss in more detail in Appendix C.
@@ 4 Experiments
For all OccamLLM results, we use Llama 3 8B Instruct [35] as the underlying language model. As such, we call our model OccamLlama. We use a 1 layer Complete OccamNet with primitives
@@
This single layer OccamNet can be invoked by the LLM several times during generation to perform complex arithmetic operations accurately. To use the trained OccamLlama for inference, we sample the highest probability function from OccamNet as described in Appendix D.3.
@@
To determine if a model output is correct, we parse all numbers in the model output and if one of them “matches” the correct answer, we determine that the result is correct. We mark each correct result as 100% accuracy and each incorrect result as 0% accuracy. For each model on each dataset, we report the mean accuracy and the standard error of the mean. To determine if a number matches the result, we first determine how many places after the decimal $d$ the number should be accurate to. If the number is an integer, we set $d$ to 2. Otherwise, we set $d$ to the number of places after the decimal in the model output, clipped between 2 and 5. Finally we state that a number “matches” the result if the number and the result differ by less than $10^{-d}$. We present further experiment details, including hyperparameters, prompts, etc. in Appendix A.
@@ 4.1 Simple Arithmetic Problems
Table 2: Accuracy on arithmetic tasks. OccamLlama uses Llama 3 8B Instruct as the underlying language model. Higher is better. Bold indicates best performance for each row.
@@
7B Chat 8B Instruct Turbo Code
@@
The results of these evaluations are shown in Table 2. More detailed results, including relative error and results for 3 and 5 digit arithmetic, are shown in Appendix A. OccamLlama has $100.0\pm 0.0$% accuracy on all tasks, missing 0 out of 9000 problems. On the other hand, we tested GPT 4o with Code Interpreter on fewer problems to save cost, and it missed 3 out of the 1800 problems it faced, achieving an accuracy of $99.8\pm 0.1$%.
@@
Figure 3: Accuracy of OccamLlama and baselines on mathematical problem solving tasks. Higher is better. OccamLlama achieves accuracy comparable to Llama 3 8B on benchmarks with simple arithmetic, higher accuracy than GPT 4o and GPT 4o + Code on on tasks with challenging arithmetic, and accuracy above Llama 3 8B and similar to GPT 3.5 Turbo on average.
@@ 4.2 Mathematical Problem Solving
To test the performance of OccamLlama on more general mathematical problem solving tasks, we evaluate our method and baselines on the following six benchmarks: AddSub [39], GSM8K [40], MultiArith [33], MATH401 [8], Single Eq [41], and SVAMP [42]. All but MATH401 are word problems requiring longer generation and a mix of reasoning and arithmetic capabilities. MATH401 also includes multistep arithmetic problems which require more than one call to OccamLlama.
@@
Because many of the arithmetic operations required in these datasets are relatively simple, we also create MultiArith Float, a modification of MultiArith in which we select problems which are arithmetically more challenging, while requiring similar levels of reasoning. To this end, we select prompts having input numbers that can be replaced with floats. For instance, $3.5$ feet or \$$\$39.95$ are reasonable but $3.5$ people is not. Furthermore, we sample input values from ranges larger than those appearing in the MultiArith dataset, in cases where it is reasonable. Float operations and larger additions and multiplications are more difficult for the baseline LLMs but do not make a difference for OccamLLM, so this dataset is particularly useful to show the advantages of the system we propose.
@@
Figure 3 shows the results of these evaluations. More detailed results are shown in Appendix A. On MultiArith Float and MATH401, two datasets requiring challenging arithmetic, OccamLlama outperforms not only Llama 3 8B but also GPT 4o and GPT 4o + Code. At the same time, most other datasets in this benchmark do not involve challenging arithmetic, meaning that Llama 3 8B is well suited to solve these tasks without assistance; most of the difficulty of these tasks lies in the reasoning rather than in the arithmetic computations. This is further supported by the fact that GPT 4o with Code Interpreter never substantially outperforms and sometimes underperforms GPT 4o. As such, it is remarkable that OccamLlama can achieve comparable accuracy to Llama 3 8B even when it is trained on very different data and evaluated on tasks without challenging arithmetic. The only dataset for which OccamLlama performs substantially worse than Llama 3 8B is GSM8K, but we attribute this to an imperfect OccamLlama switch, likely stemming from text which is outside of the switch training distribution (see Section 4.3).
@@
In Figure 4, we show example generations from OccamLlama for both arithmetic and reasoning tasks. These generations demonstrate how the OccamLlama switch learns to balance OccamNet outputs with LLM outputs, effectively distributing the work between a reasoner (Llama) and a calculator (OccamNet). Because the language model is unaware of the OccamLlama system, its generations behave as if it possesses an interior calculator even though it is actually using a tool. In this way, we combine the benefits of a language model finetuned on arithmetic with the benefits of a language model finetuned to use code for arithmetic, all without any finetuning.
@@
Figure 4: Examples from Llama 3 8B Instruct and OccamLlama on (top) an arithmetic problem and (bottom) a mathematical reasoning problem from the MultiArith Float dataset. In OccamLlama, the LLM performs reasoning, the switch predicts when to use OccamNet, and OccamNet performs arithmetic operations. OccamNet’s inputs and outputs are highlighted in purple and green, respectively.
@@ 4.3 Limitations
In our experiments, we use a single-layer OccamNet as the symbolic network, enabling evaluation of single-operation arithmetic problems. This sometimes poses a challenge on reasoning problems, when the base language model generates compound expressions containing more than one operation, such as $3+5+7=$. A single-layer OccamNet cannot evaluate these expressions. We attempt to overcome this by prompting Llama to break down compound expressions into multiple steps, but we find it difficult to coerce Llama to follow these instructions. Another challenge is that Llama often generates expressions in fractions or percentages, which also constitute compound expressions that are not properly handled by the OccamLLM system. Fortunately, we observed that these compound expressions were typically simple enough for the LLM to evaluate without OccamNet. Therefore, in our experiments, we trained the OccamLLM switch to avoid using OccamNet for compound operations. In future work, we can explore other solutions such as integrating a two-layer OccamNet as the symbolic network. We found that these issues are particularly acute in the GSM8K [40] dataset, where the expressions generated by Llama are not prevalent in the switch training data, causing it to sometimes incorrectly trigger OccamNet and degrade performance.
@@ 5 Discussion
We presented OccamLLM, a system enabling exact and interpretable language model arithmetic in a single autoregressive step. Our method does not require modifying the weights of the underlying language model, thereby avoiding risks of catastrophic forgetting. Furthermore, our method avoids security risks arising from running code generated by a language model.

### v4

@@ Abstract
Despite significant advancements in text generation and reasoning, Large Language Models (LLMs) still face challenges in accurately performing complex arithmetic operations. Language model systems often enable LLMs to generate code for arithmetic operations to achieve accurate calculations. However, this approach compromises speed and security, and fine-tuning risks the language model losing prior capabilities. We propose a framework that enables exact arithmetic in a single autoregressive step, providing faster, more secure, and more interpretable LLM systems with arithmetic capabilities. We use the hidden states of a LLM to control a symbolic architecture that performs arithmetic. Our implementation using Llama 3 with OccamNet as a symbolic model (OccamLlama) achieves 100% accuracy on single arithmetic operations (+,,×,÷$+,-,\times,\div,\sin{},\cos{},\log{},\exp{},\sqrt{}$), outperforming GPT 4o with and without a code interpreter. Furthermore, OccamLlama outperforms GPT 4o with and without a code interpreter on average across a range of mathematical problem solving benchmarks, demonstrating that OccamLLMs can excel in arithmetic tasks, even surpassing much larger models. We will make our code public shortly.
@@ 1 Introduction
    We benchmark on mathematical problem solving tasks, showing that OccamLlama can sustain long generations. OccamLlama outperforms both GPT 4o and GPT 4o with code interpreter on average across the benchmarks we tested.
@@ 2 Related Work
Table 1: OccamLLM is the only approach to improving the arithmetic capabilities of a pretrained LLM which 1) enables single-pass arithmetic, 2) does not risk catastrophic forgetting from finetuning, 3) does not require arbitrary code execution, and 4) provides an interpretable process.
@@ 3.1 OccamLLM: Combining a Language Model with a Symbolic Model
In the subsections below, we describe the OccamLLM system which from our experiments we find to be most performant, even oupterforming GPT 4o in several benchmarks. For an analysis of alternate architectures and losses, see Appendix D.
@@ 3.1.1 OccamNet
OccamNet is a symbolic architecture that provides an interpretable way of parametrizing probability distributions over a space of functions [11]. We leave a more thorough explanation of OccamNet to [11] and Appendix E, describing only the relevant components here.
@@ 3.1.2 OccamLLM Decoder
where $\mathbf{h}_{j}$ are the hidden states of the $j$th layer of the language model, $w_{i,j}$ are trainable weights, $\text{MLP}_{i}$ are two-layer multilayer perceptrons (MLPs), and $\mathbf{W}^{*(i)}$ are untrained weights which initialize all functions to have approximately equal probabilities according to the initialization scheme described in [11] and explained in Appendix E.4.
@@ 3.2.1 OccamNet Decoder Training Data
For training the decoder that controls the weights of OccamNet, we created two types of examples, single queries and concatenated queries. For single queries, we select a single prompt from the problems generated as discussed in Section 3.2. We use the Llama 3 Instruct chat template and fill in the query as the user input and the result as the assistant response, prepending “Answer = ” to the later in randomly selected samples (see Appendix A.1.1 for further details). For the concatenated queries of examples, we select a random number of prompts and concatenate the query-response pairs without using the Llama 3 Instruct chat template. The OccamNet decoder is trained to predict only the results of the last query in the sequence. This strategy helps OccamLLM to learn which operation to perform without becoming confused by earlier text, which is useful for continuous generation. To create the training dataset, each example is sampled by first randomly selecting whether to create a single or concatenated query, then randomly selecting the type(s) of prompt(s) used, and finally randomly sampling the input values from the range corresponding to each selected prompt.
@@ 3.2.2 OccamLLM Switch Training Data
To train the switch, we generate examples of possible LLM outputs for given input expressions and label the outputs with sequences of 0s or 1s corresponding to whether the language model or the OccamNet output should be used for the next token. Some examples correspond to the prompts described in Section 3.2. For such examples, the LLM output is set to “The answer is” or “Answer = ” and the label sequence is all 0s with a 1 at the last token to indicate the system should use OccamNet only to compute the answer. We also manually created and labeled several other examples for diverse scenarios to explicitly teach the system in which cases it should or should not use OccamNet (see Appendix A.1.2 for further details).
To create the training dataset, we concatenate a random number of the above user input - assistant output pairs in a conversational fashion, using the Llama 3 Instruct chat template.
@@ 3.3 OccamLLM Training
We train the OccamLLM decoder and the switch separately, as they do not share weights. In all cases, the weights of the LLM are kept frozen. In the first step, we train the system to predict the answer to examples generated by the method explained in Section 3.2.1. The OccamNet decoder processes the hidden states corresponding to the last token of the response and sets the weights of OccamNet such that the correct arithmetic expression is sampled. In this step, we use a rescaled REINFORCE [34] loss, which can also be interpreted as a Monte-Carlo estimate of the cross-entropy loss (see Appendix D.2):
@@
Minimizing this loss steers the decoder towards assigning higher probabilities to the functions that maximize the reward $R(f(x),y)$, which measures the similarity between the correct answer $y$ and the prediction of OccamNet $f(x)$. We find setting $R(f(x),y)=1$ if $f(x)=y$, and 0 otherwise, most effective. We discuss the OccamNet loss in more detail in Appendix D.
@@ 4 Experiments
For all OccamLLM results, we use Llama 3 8B Instruct and Llama 3 70B Instruct [35] as the underlying language models. As such, we call our models OccamLlama 8B and OccamLlama 70B, respectively. We use a 1 layer Complete OccamNet with primitives
@@
This single layer OccamNet can be invoked by the LLM several times during generation to perform complex arithmetic operations accurately. To use the trained OccamLlama for inference, we sample the highest probability function from OccamNet as described in Appendix E.3.
@@
To determine if a model output is correct, we parse all numbers in the model output and if one of them “matches” the correct answer, we determine that the result is correct. We mark each correct result as 100% accuracy and each incorrect result as 0% accuracy. For each model on each dataset, we report the mean accuracy and the standard error of the mean. To determine if a number matches the result, we first determine how many places after the decimal $d$ the number should be accurate to. If the number is an integer, we set $d$ to 2. Otherwise, we set $d$ to the number of places after the decimal in the model output, clipped between 2 and 5. Finally we state that a number “matches” the result if the number and the result differ by less than $10^{-d}$. We present further experiment details, including additional experiments, hyperparameters, and prompts in Appendix A.
@@ 4.1 Simple Arithmetic Problems
Table 2: Accuracy on arithmetic tasks, in percentages. The OccamLlama column corresponds to the results of both OccamLlama 8B and OccamLlama 70B. Higher is better. Bold indicates best performance for each row.
@@
8B / 70B 7B Chat 8b Instruct Turbo Code
@@
The results of these evaluations are shown in Table 2. More detailed results, including relative error and results for 3- and 5-digit arithmetic, are shown in Appendix A.5.
Both OccamLlama 8B and 70B have $100.0\pm 0.0$% accuracy on all tasks, missing 0 out of 9000 problems. On the other hand, we tested GPT 4o with Code Interpreter on fewer problems to save cost, and it missed 3 out of the 1800 problems it faced, achieving an accuracy of $99.8\pm 0.1$%.
@@
Figure 3: Accuracy of OccamLlama and baselines on mathematical problem solving tasks. Higher is better. OccamLlama 8B achieves accuracy comparable to Llama 3 8B on benchmarks with simple arithmetic, higher accuracy than GPT 4o and GPT 4o + Code on on tasks with challenging arithmetic, and accuracy above Llama 3 8B and similar to GPT 3.5 Turbo on average. OccamLlama 70B outperforms GPT 4o and GPT 4o + Code on average.
@@ 4.2 Mathematical Problem Solving
To test the performance of OccamLlama on more general mathematical problem solving tasks, we evaluate our method and baselines on the following six benchmarks: AddSub [39], GSM8K [40], MultiArith [33], MATH401 [8], Single Eq [41], and SVAMP [42]. All but MATH401 are word problems requiring longer generation and a mix of reasoning and arithmetic capabilities. MATH401 also includes multistep arithmetic problems which require more than one call to OccamLlama. We selected these datasets (including the MultiArith Float dataset described below) before testing any methods on them to ensure unbiased selection of benchmarks.
@@
Because many of the arithmetic operations required in these datasets are relatively simple, we also create MultiArith Float, a modification of MultiArith in which we select problems which are arithmetically more challenging, while requiring similar levels of reasoning. To this end, we select prompts having input numbers that can be replaced with floats. For instance, $3.5$ feet or \$$\$39.95$ are reasonable but $3.5$ people is not. Furthermore, we sample input values from ranges larger than those appearing in the MultiArith dataset, in cases where it is reasonable. Float operations and larger additions and multiplications are more difficult for the baseline LLMs but do not make a difference for OccamLLM, so this dataset is particularly useful to show the advantages of the system we propose. Figure 3 shows the results of these evaluations. More detailed results are shown in Appendix A.5.
OccamLlama 70B outperforms both GPT 4o and GPT 4o + Code on average across the benchmarks, demonstrating OccamLlama’s strong mathematical problem solving capability. We also note that GPT 4o + Code does not outperform GPT 4o on average, suggesting that existing implementations of LLMs with code generation may not help with mathematical problem solving.
@@
We now consider the performance of OccamLlama 8B, the smaller OccamLlama model. On MultiArith Float and MATH401, two datasets requiring challenging arithmetic, OccamLlama 8B outperforms not only Llama 3 8B but also GPT 4o and GPT 4o + Code. At the same time, most other datasets in this benchmark do not involve challenging arithmetic, meaning that Llama 3 8B is well suited to solve these tasks without assistance; most of the difficulty of these tasks lies in the reasoning rather than in the arithmetic computations. This is further supported by the fact that GPT 4o with Code Interpreter never substantially outperforms and sometimes underperforms GPT 4o on these tasks. As such, it is remarkable that OccamLlama 8B can achieve comparable accuracy to Llama 3 8B even when it is trained on very different data and evaluated on tasks without challenging arithmetic.
The only datasets for which OccamLlama 8B performs noticeably worse than Llama 3 8B are GSM8K and Single Eq, but we believe this results from an imperfect OccamLlama switch, likely stemming from text which is outside of the switch training distribution (see Section 4.3). Fortunately, in Appendix C, we find that the OccamNet decoder is quite robust to out of distribution data and that both the OccamNet and switch decoders generalize well to unseen languages. This suggests that, with relatively little data, it should be possible to teach the switch to handle these unseen cases, something we leave for future work.
@@
In Figure 4, we show example generations from OccamLlama 8B for both arithmetic and reasoning tasks. These generations demonstrate how the OccamLlama switch learns to balance OccamNet outputs with LLM outputs, effectively distributing the work between a reasoner (Llama) and a calculator (OccamNet). Because the language model is unaware of the OccamLlama system, its generations behave as if it possesses an interior calculator even though it is actually using a tool. In this way, we combine the benefits of a language model finetuned on arithmetic with the benefits of a language model finetuned to use code for arithmetic, all without any finetuning.
@@
Figure 4: Examples from Llama 3 8B Instruct and OccamLlama 8B on (top) an arithmetic problem and (bottom) a mathematical reasoning problem from the MultiArith Float dataset. In OccamLlama, the LLM performs reasoning, the switch predicts when to use OccamNet, and OccamNet performs arithmetic operations. OccamNet’s inputs and outputs are highlighted in purple and green, respectively.
@@ 4.3 Limitations
In our experiments, we use a single-layer OccamNet as the symbolic network, enabling evaluation of single-operation arithmetic problems. This sometimes poses a challenge on reasoning problems when the base language model generates compound expressions requiring more than one operation to evaluate, such as $3+5+7=$. A single-layer OccamNet cannot evaluate these expressions. We attempted to overcome this by prompting Llama to break down compound expressions into multiple steps, but we find it difficult to coerce Llama to follow these instructions. Another challenge is that Llama often generates expressions in fractions or percentages, which also constitute compound expressions that are not properly handled by the OccamLLM system. Fortunately, we observed that these compound expressions were typically simple enough for the LLM to evaluate without OccamNet. Therefore, in our experiments, we trained the OccamLLM switch to avoid using OccamNet for compound operations, largely mitigating this issue. Future work could explore other solutions such as integrating a two-layer OccamNet as the symbolic network. We found that these issues are particularly acute in the GSM8K and Single Eq datasets, where the expressions generated by Llama are not prevalent in the switch training data, causing it to sometimes incorrectly trigger OccamNet and degrade performance, as discussed more in Appendix A.5.
@@ 5 Discussion
We presented OccamLLM, a system enabling exact and interpretable language model arithmetic in a single autoregressive step. Our method does not require modifying the weights of the underlying language model, thereby avoiding risks of catastrophic forgetting. Furthermore, our method avoids security risks arising from running code generated by a language model while outperforming top LLM code generation methods (GPT 4o + Code) on average across our benchmarks.
We benchmarked our method on challenging arithmetic tasks, achieving 100% accuracy where GPT 4o achieves only 40% performance on average. We also benchmarked our method on mathematical problem solving tasks, demonstrating that the OccamLlama switch can accurately balance the LLM for reasoning and OccamNet for arithmetic, outperforming even GPT 4o and GPT 4o with Code Interpreter on average.
Our work could enable smaller LLMs to be as performant as much larger LLMs in arithmetic. Moreover, integrating OccamLLM with larger LLMs like GPT 4o could further improve their arithmetic abilities without requiring a code interpreter. Furthermore, at present, OccamLLM may not integrate with more advanced decoding techniques such as speculative decoding [38, 43]. We hope to explore these avenues in future work.
6 Broader Impact
We believe that, in addition to enabling fast, safe, and interpretable arithmetic, OccamLLM demonstrates a new paradigm for tool use. As a proof of concept for more complex tool use, we further train OccamLlama 8B with a two layer Complete OccamNet with the primitives